Calls originate from Dryden according to a rate 12 Poisson process. 3/4 are local and 1/4 are long distance. Local calls last an average of 10 minutes, while long distance calls last an average of 5 minutes. Let M be the number of local calls and N the number of long distance calls in equilibrium. Find the distribution of (M,N). what is the number of people on the line.
Answer #1

here for an hour expected number of Local calls =12*0.75 =9

therefore expected amount of time spent by Local calls in 8 hours=9*8*10=720

for an hour expected number of Long distance calls =12*0.25 =3

therefore expected amount of time spent by Long distance calls in 8 hours=3*8*5=120

therefore

expected number of minutes that all callers from Dryden will spend on the phone during an eight hour period =720+120=840